# thanks: $0$-0002-5209-872Xthanks: $0$-0002-0428-8430

### v1

@@ Abstract
A nearby supernova will carry an unprecedented wealth of information about astrophysics, nuclear physics, and particle physics. Because supernova are fundamentally neutrino driven phenomenon, our knowledge about neutrinos – particles that remain quite elusive – will increase dramatically with such a detection. One of the biggest open questions in particle physics is related to the masses of neutrinos. Here we show how a galactic supernova provides information about the masses of each of the three mass eigenstates individually, at some precision, and is well probed at JUNO. This information comes from several effects including time delay and the physics within the supernova. The time delay feature is strongest during a sharp change in the flux such as the neutronization burst; additional information may also come from a QCD phase transition in the supernova or if the supernova forms a black hole. We consider both standard cases as dictated by local oscillation experiments as well as new physics motivated scenarios where neutrino masses may differ across the galaxy.
@@ I Introduction
Another method of probing the absolute neutrino mass scale is via measurements of neutrinos from a galactic supernova (SN) Zatsepin (1968), see also Loredo and Lamb (2002); Nardi and Zuluaga (2004, 2005); Pagliaroli et al. (2010); Lu et al. (2015); Abe et al. (2018); Hansen et al. (2020); Pompa et al. (2022); Pitik et al. (2022); Brdar and Xu (2022), which is the focus of this paper. Existing studies typically assume equal masses for each neutrino, incompatible with oscillation results. We develop the theory here, for the first time, for the impact of three separate mass states on both the propagation of neutrinos from a SN to the Earth, but also the impact within the SN. We consider the JUNO detector Abusleme et al. (2024), under construction now, as the target due to its large volume and low thresholds. Finally, we consider three possible sharp features with which to observe the time-delay signature and thus extract the masses of neutrinos: the neutronization burst Mirizzi et al. (2016) which is quite likely to exist, a QCD phase transition inside the SN Sagert et al. (2009); Fischer et al. (2011, 2018); Zha et al. (2021); Fischer (2021); Kuroda et al. (2022); Bauswein et al. (2022); Lin et al. (2024); Pitik et al. (2022) which may exist, and finally the possibility that the SN forms a black hole (BH) Sekiguchi and Shibata (2011); Gullin et al. (2022) which happens for some subset of SN explosions.
Figure 1: A schematic showing the effects in play from the SN explosion, propagation, and detection. Different masses will potentially affect the adiabatic flavor change in the SN, they will affect the propagation as a function of neutrino energy, and then each mass stage is projected back to a given flavor state at the Earth.
@@ II Possibility of Spatially Evolving Neutrino Masses
In Davoudiasl et al. (2018) they proposed that neutrino masses are exactly zero except near DM due to a long-range mediator. Thus the measured parameters would be related to the local DM density and the neutrino masses would be higher in the galactic center proportional to the DM density there and the individual masses would remain proportional to those near the Earth, but the neutrino mass separations would become much larger. We note here that the additional presence of a bare mass term arising from any typical scenario (e.g. seesaw) with possibly a different sign, could lead to a wide range of non-trivial phenomenology, masses, and mass differences.
@@ III.1 Neutrino Mixing
This allows us to create a single framework that covers not only both the normal mass ordering ($L=1$, $M=2$, and $H=3$) and the inverted mass ordering ($L=3$, $M=1$, $H=2$), but any other possible configuration.
We assume (and will soon show) that neutrinos are produced with an energy and in a density larger than all resonances for any available $\Delta m^{2}$. Thus the neutrinos will evolve through both level crossings. Under these set of reasonable and justified assumption, in general, we find
@@
While these expressions (eqs. 11-14) look similar to those above (eqs. 1-8), the careful placement of the mass orderings allows for a completely general description of neutrino masses and is necessary for a time delay calculation involving each neutrino mass state separately both for assumed true values of the neutrino masses, as well as alternate test values when computing statistical significances. These equations represent an important result of this work.
@@ III.2.2 Three-Flavor Picture
We are in a three flavor picture, however, which complicates things somewhat. Since the formulas are derived in a two-flavor picture and involve an integrand of the form
@@
which is the the typical matter correction factor, we know how to best adjust these for a three flavor picture, see Denton and Parke (2019), where $a=2EV_{\rm CC}$ quantifies the matter effect.
@@
For arbitrary $m_{i}$, these two flavor approximations may not hold, and thus the corresponding angles may not be correct, but they should set the approximate scale which we will show is safe by many orders of magnitude.
@@ III.2.3 Numerical Studies
We show the regions of $P_{j}\geq 0.1$ and $P_{j}\geq 0.9$ for different values of the two relevant masses for each relevant angle: $\theta_{12}$ and $\theta_{13}$, along with the zoomed in region, at $E=10$ MeV, in fig. 2. The same study is repeated at $E=100$ MeV in fig. 3, although this is just included to show the energy scaling; for the time delay calculation, the contribution to the information comes dominantly for lower energy neutrinos.
We see that, except for a tiny sliver of parameter space, the effect is completely negligible. It only begins to become relevant for energies higher than are relevant for the time delay effect and at masses that are very small and much closer together than anticipated from the $\Delta m^{2}$’s. A new physics scenario involving masses this close together in the galactic center that then returns to the measured oscillation parameters in our solar system would be quite finely tuned. In addition, the only region where a physical $\Delta m^{2}$ begins to become close to a noticeable jump probability is for masses much larger than consistent with cosmology.
@@ IV Time Delay Features
where we note that the mass state that should be applied for a given time delay depends on the prefactor of the PMNS matrix that is applied. This makes it manifest why it is essential to not apply unitarity conditions when writing the flavor transformations.
Having a time delay effect does not necessarily lead to an easily identifiable signal without a sharp feature in the spectrum. That is, ideally the spectrum will suddenly and dramatically change due to some large scale feature of the explosion. We explore several such possibilities throughout this section and perform numerical studies for three of them in section VI.
@@ IV.2 QCD Phase Transition
A QCD phase transition, i.e. a quark hadron phase transition, may occur in SN. If so, the SN explosion includes a first order phase transition from nuclear matter to quark matter. This can produce a new source of energy to drive a second shock wave restarting a stalled shock. See Sagert et al. (2009); Fischer et al. (2011, 2018); Zha et al. (2021); Fischer (2021); Kuroda et al. (2022); Bauswein et al. (2022); Jakobus et al. (2022); Lin et al. (2024); Pitik et al. (2022) for various discussions of this phenomenon.
@@
Figure 5: The expected event rates for a SN at 10 kpc detected by JUNO as a function of energy and time for three different neutrino mass configurations. In the top left panel the neutrinos are massless, in the top right panel the neutrinos are in a varied high-mass scenario (see subsection VI.1 and table 1), and in the bottom panel the neutrinos are all quite heavy. Note that in the top right panel, some of the low-energy events are somewhat delayed due to $m_{3}$ large, but the effect of the remaining events is smaller. Additional plots for neutronization burst and black hole formation are provided in Appendix A.
@@ IV.3 Black Hole Formation
We parameterize BH formation by rescaling the flux, mean neutrino energy, and pinching parameter via a sigmoid function. This ensures that the evolution of these quantities during BH formation closely matches the numerical results presented in Gullin et al. (2022) and includes the realistic turn off time for the neutrino flux which has a characteristic time $\mathcal{O}(2)$ ms. As with the QCD phase transition, we include a window of width 40 ms around the BH formation time.
@@ IV.4 SASI
Another possible feature that gives rise to a time dependent signature is the Standing Accretion Shock Instability (SASI) Blondin et al. (2003); Foglizzo et al. (2015); Lin et al. (2020). This leads to an oscillatory signature in the neutrino flux. When combined with time delay, which depends on energy, it could be possible to extract information on neutrino masses. We do not consider this effect in this study because the magnitude of this effect varies considerably from simulation to simulation, the effect tends to get largely washed out if the progenitor carries moderate or large angular momentum, and marginalizing over the uncertainties in the oscillatory signature will likely wash out the effect considerably. Nonetheless, if a galactic SN neutrino signal is detected and if it carries evidence of SASI, a time delay analysis on that region of the data should also be performed.
@@ V.1 JUNO
In this article we will focus on detection with JUNO Abusleme et al. (2022, 2024). Even though the detector volume of JUNO will be smaller than some other large-volume neutrino detectors, it is expected to come online before other next generation experiments DUNE and HK. In addition, due to the usage of liquid scintillator and state-of-the-art photomultiplier tubes, large overburden, and good radiopurity, JUNO is expected to achieve fairly low thresholds. Since the time delay effect scales like $E^{-2}$, probing lower energy neutrinos results in a greater advantage for determining neutrino masses. In addition, this will reduce the impact due to the uncertainty in the nature of the SN neutrino flux itself by measuring the time delay effect across a range of energies. We implement JUNO’s detection capabilities as described in SNowGLoBES Scholberg et al. (2021) including cross sections and experimental details. We also find that JUNO will likely have excellent timing resolution $\lesssim 1$ ns, better than needed for typical SN, given the $<10$ cm spatial resolution Qian et al. (2021).
@@ VI.1 Benchmarks
Cosmological measurements constrain the quantity $\sum_{i}m_{i}$. The data indicates that this quantity is consistent with zero (actually slightly preferring negative values) Aghanim et al. (2020); Adame et al. (2024); Craig et al. (2024); Jiang et al. (2024) and is putting pressure on the inverted mass ordering (the case where the neutrino that is least $\nu_{e}$ is the lightest), which has a larger minimum sum of the neutrino masses. In fact, the data is also somewhat disfavoring the normal ordering as well, at low significance. While future data is expected to have good sensitivity to measuring this quantity, the current situation with the data, as well as tensions in cosmology Abdalla et al. (2022), is somewhat complicated. In light of this, we use the constraint from Planck alone (TT, TE, EE, lowE, and lensing) as one of our benchmarks $\sum_{i}m_{i}<0.24$ eV Aghanim et al. (2020).
A cleaner but less competitive measurement comes from KATRIN looking at the endpoint of the tritium spectrum Aker et al. (2024) and is expected to gain another factor of $\sim 2$ in precision, although this will still not be enough to be competitive with the cosmological constraint. Other probes are possible as well such as neutrinoless double beta decay provided that neutrinos have Majorana mass terms and the light neutrino exchange model is valid.
Combining this, our benchmark points, shown in table 1, are either the minimum masses allowed by oscillations, or the maximum masses allowed by other data sets. We also include one high mass scenario that does not take into account the measured $\Delta m^{2}_{ij}$’s, that could be due to new physics. We have also investigated other scenarios such as the upper limit on the masses from various cosmological data sets including CMB, BAO, and the latest BAO updates from DES and DESI. We found that the constraints in those cases were quite similar to the lowest possible masses.
@@
Figure 6: The expected preferred regions given a detection of a SN at 10 kpc by JUNO under four different hypotheses about the true neutrino masses (see table 1). Note that in this case, only information from the QCD phase transition is used to determine the neutrino masses. The contours are drawn at 1, 2, and 3 $\sigma$ assuming 2 dofs and Wilks’ theorem, and the shaded regions corresponds to the 1 $\sigma$ region for 1 dof. The pluses and vertical dashed lines represent the true values.
@@ VII Discussion
The numerical studies have yielded many interesting and non-trivial results. Perhaps the most compelling result found is the fact that in general there is information that can distinguish among the three different mass states in a galactic SN. That is, depending on the true masses of neutrinos, each individual mass state can be uniquely measured. The uniqueness statement is based on the fact that we assume that the PMNS matrix is determined. Thus it is not surprising that the constraints on $m_{3}$ tend to be poorer because $|U_{e3}|$ is much smaller than $|U_{e1}|$ and $|U_{e2}|$. In fact, we believe that any estimate of neutrino masses from a galactic SN should use the full description including the explicit contribution from each separate mass term as described in eqs. 11-14.

### v2

@@ Abstract
A nearby supernova will carry an unprecedented wealth of information about astrophysics, nuclear physics, and particle physics. Because supernova are fundamentally neutrino driven phenomenon, our knowledge about neutrinos – particles that remain quite elusive – will increase dramatically with such a detection. One of the biggest open questions in particle physics is related to the masses of neutrinos. Here we show how a galactic supernova provides information about the masses of each of the three mass eigenstates individually, at some precision, and is well probed at JUNO. This information comes from several effects including time delay and the MSW effect within the supernova. The time delay feature is strongest during a sharp change in the flux such as the neutronization burst; additional information may also come from a QCD phase transition in the supernova or if the supernova forms a black hole. We consider both standard cases as dictated by local oscillation experiments as well as new physics motivated scenarios where neutrino masses may differ across the galaxy.
@@ I Introduction
Another method of probing the absolute neutrino mass scale is via measurements of neutrinos from a galactic supernova (SN) Zatsepin (1968), see also Loredo and Lamb (2002); Nardi and Zuluaga (2004, 2005); Pagliaroli et al. (2010); Lu et al. (2015); Abe et al. (2018); Hansen et al. (2020); Pompa et al. (2022); Pitik et al. (2022); Brdar and Xu (2022); Parker and Wurm (2024), which is the focus of this paper. Existing studies typically assume equal masses for each neutrino, incompatible with oscillation results. We develop the theory here, for the first time, for the impact of three separate mass states on both the propagation of neutrinos from a SN to the Earth, and also the impact within the SN. We consider the JUNO detector Abusleme et al. (2024), under construction now, as the target due to its large volume and low thresholds. Finally, we consider three possible sharp features with which to observe the time-delay signature and thus extract the masses of neutrinos: the neutronization burst Mirizzi et al. (2016) which is quite likely to exist, a QCD phase transition inside the SN Sagert et al. (2009); Fischer et al. (2011, 2018); Zha et al. (2021); Fischer (2021); Kuroda et al. (2022); Bauswein et al. (2022); Lin et al. (2024); Pitik et al. (2022) which may exist, and finally the possibility that the SN forms a black hole (BH) Sekiguchi and Shibata (2011); Gullin et al. (2022) which happens for some subset of SN explosions.
Figure 1: A schematic showing the effects in play from the SN explosion, propagation, and detection. Different masses will potentially affect the adiabatic flavor change in the SN; they will affect the propagation as a function of neutrino energy, and then each mass stage is projected back to a given flavor state at the Earth.
@@ II Possibility of Spatially Evolving Neutrino Masses
In Davoudiasl et al. (2018) they proposed that neutrino masses are exactly zero except near DM due to a long-range mediator. Thus the measured parameters would be related to the local DM density and the neutrino masses would be higher in the galactic center proportional to the DM density there and the individual masses would remain proportional to those near the Earth, but the neutrino mass separations would become much larger. We note here that the additional presence of a bare mass term arising from any typical scenario (e.g. seesaw) with possibly a different sign, could lead to a wide range of nontrivial phenomenology, masses, and mass differences. For example, it could lead to mass orderings completely different from the two standard ones: the normal ordering $m_{1}<m_{2}<m_{3}$ and the inverted ordering $m_{3}<m_{1}<m_{2}$. In fact, any of the six possible mass orderings can be realized in these new physics scenarios in the galactic center and still be consistent with oscillation data on the Earth.
@@ III.1 Neutrino Mixing
This allows us to create a single framework that covers not only both the normal mass ordering ($L=1$, $M=2$, and $H=3$) and the inverted mass ordering ($L=3$, $M=1$, $H=2$), but also any other possible configuration.
We assume (and will soon show) that neutrinos are produced with an energy and in a density larger than all resonances for any available $\Delta m^{2}$. Thus the neutrinos will evolve through both level crossings. Under this set of reasonable and justified assumptions, in general, we find
@@
While these expressions (eqs. 11-14) look similar to those above (eqs. 1-8), the careful placement of the mass orderings allows for a completely general description of neutrino masses and is necessary for a time delay calculation involving each neutrino mass state separately both for assumed true values of the neutrino masses, as well as alternate test values when computing statistical significances. These equations enable the calculation of the neutrino flux at Earth for any mass ordering and represent an important result of this work.
@@ III.2.2 Three-Flavor Picture
We are in a three-flavor picture, however, which complicates things somewhat. Since the formulas are derived in a two-flavor picture and involve an integrand of the form
@@
which is the the typical matter correction factor, we know how to best adjust these for a three-flavor picture, see Denton and Parke (2019), where $a=2EV_{\rm CC}$ quantifies the matter effect.
@@
For arbitrary $m_{i}$, these two-flavor approximations may not hold, and thus the corresponding angles may not be correct, but they should set the approximate scale which we will show is safe by many orders of magnitude.
@@ III.2.3 Numerical Studies
We show the regions of $P_{j}\geq 0.1$ and $P_{j}\geq 0.9$ for different values of the two relevant masses for each relevant angle: $\theta_{12}$ and $\theta_{13}$, along with the zoomed-in region, at $E=10$ MeV, in fig. 2. The same study is repeated at $E=100$ MeV in fig. 3, although this is just included to show the energy scaling; for the time delay calculation, the contribution to the information comes dominantly for lower-energy neutrinos.
We see that, except for a tiny sliver of parameter space, the effect is completely negligible. It only begins to become relevant for energies higher than are relevant for the time delay effect and at masses that are very small and much closer together than anticipated from the $\Delta m^{2}$’s. A new physics scenario involving masses this close together in the galactic center that then returns to the measured oscillation parameters in our solar system would be quite finely tuned. In addition, the only region where a physical $\Delta m^{2}$, indicated by the orange lines, begins to become close to a noticeable jump probability is for masses much larger than consistent with cosmology or KATRIN.
@@ IV Time Delay Features
where we note that the mass state that should be applied for a given time delay depends on the prefactor of the PMNS matrix that is applied. This makes it manifest why it is essential to not apply unitarity conditions to rewrite the flavor transformations into simpler looking expressions.
Having a time delay effect does not necessarily lead to an easily identifiable signal without a sharp feature in the spectrum. That is, ideally the spectrum will suddenly and dramatically change due to some large-scale feature of the explosion. We explore several such possibilities throughout this section and perform numerical studies for three of them in section VI. We note that while the exact details of these features may have sizable uncertainties in the magnitude and shape of the effect, they are unlikely to be degenerate with the time delay signature: a shifting of the low energy events into later times.
@@ IV.2 QCD Phase Transition
A QCD phase transition, i.e., a quark hadron phase transition, may occur in SN. If so, the SN explosion includes a first-order phase transition from nuclear matter to quark matter. This can produce a new source of energy to drive a second shock wave restarting a stalled shock. See Sagert et al. (2009); Fischer et al. (2011, 2018); Zha et al. (2021); Fischer (2021); Kuroda et al. (2022); Bauswein et al. (2022); Jakobus et al. (2022); Lin et al. (2024); Pitik et al. (2022) for various discussions of this phenomenon.
@@
Figure 5: The expected event rates for a SN at 10 kpc detected by JUNO as a function of energy and time for three different neutrino mass configurations around the QCD phase transition. In the top left panel the neutrinos are massless, in the top right panel the neutrinos are in a varied high-mass scenario (see subsection VI.1 and table 1), and in the bottom panel the neutrinos are all quite heavy. Note that in the top right panel, some of the low-energy events are somewhat delayed due to $m_{3}$ large, but the effect of the remaining events is smaller. Additional plots for neutronization burst and black hole formation are provided in Appendix A.
@@ IV.3 Black Hole Formation
We parameterize BH formation by rescaling the flux, mean neutrino energy, and pinching parameter via a sigmoid function. This ensures that the evolution of these quantities during BH formation closely matches the numerical results presented in Gullin et al. (2022) and includes the realistic turn-off time for the neutrino flux which has a characteristic time $\mathcal{O}(2)$ ms. As with the QCD phase transition, we include a window of width 40 ms around the BH formation time.
@@ IV.4 SASI
Another possible feature that gives rise to a time-dependent signature is the Standing Accretion Shock Instability (SASI) Blondin et al. (2003); Foglizzo et al. (2015); Lin et al. (2020). This leads to an oscillatory signature in the neutrino flux. When combined with time delay, which depends on energy, it could be possible to extract information on neutrino masses. We do not consider this effect in this study because the magnitude of this effect varies considerably from simulation to simulation, the effect tends to get largely washed out if the progenitor carries moderate or large angular momentum, and marginalizing over the uncertainties in the oscillatory signature will likely wash out the effect considerably. Nonetheless, if a galactic SN neutrino signal is detected and if it carries evidence of SASI, a time delay analysis on that region of the data should also be performed.
@@ V.1 JUNO
In this article, we will focus on detection with JUNO Abusleme et al. (2022, 2024). Even though the detector volume of JUNO will be smaller than some other large-volume neutrino detectors, it is expected to come online before other next-generation experiments DUNE and HK. In addition, due to the usage of liquid scintillator and state-of-the-art photomultiplier tubes, large overburden, and good radiopurity, JUNO is expected to achieve fairly low thresholds. Since the time delay effect scales like $E^{-2}$, probing lower energy neutrinos results in a greater advantage for determining neutrino masses. In addition, this will reduce the impact due to the uncertainty in the nature of the SN neutrino flux itself by measuring the time delay effect across a range of energies where the dominant contribution to the event rate comes in the energy range $5\lesssim E/$MeV50less-than-or-similar-toabsent50$\lesssim$ 50≲ 50. We implement JUNO’s detection capabilities as described in SNowGLoBES Scholberg et al. (2021) including cross sections and experimental details which also accounts for neutrino energy smearing for all detection channels. We also find that JUNO will likely have excellent timing resolution $\lesssim 1$ ns, better than needed for typical SN, given the $<10$ cm spatial resolution Qian et al. (2021).
@@ VI.1 Benchmarks
Cosmological measurements constrain the quantity $\sum_{i}m_{i}$. The data indicates that this quantity is consistent with zero (actually slightly preferring negative values) Aghanim et al. (2020); Adame et al. (2024); Craig et al. (2024); Jiang et al. (2024) and is putting pressure on the inverted mass ordering (the case where the neutrino that is least $\nu_{e}$ is the lightest), which has a larger minimum sum of the neutrino masses. In fact, the data is also somewhat disfavoring the normal ordering as well, at low significance. While future data is expected to have good sensitivity to measuring this quantity, the current situation with the data, as well as tensions in cosmology Abdalla et al. (2022); Poudou et al. (2025), is somewhat complicated. In light of this, we use the constraint from Planck alone (TT, TE, EE, lowE, and lensing) as one of our benchmarks $\sum_{i}m_{i}<0.24$ eV Aghanim et al. (2020).
A cleaner but less competitive measurement comes from KATRIN looking at the endpoint of the tritium spectrum Aker et al. (2024) [^5] See also Project 8 Esfahani et al. (2022) as well as the electron capture proposals HOLMES Becker et al. (2019) and ECHo Gastaldo et al. (2014). and is expected to gain another factor of $\sim 2$ in precision, although this will still not be enough to be competitive with the cosmological constraint. Other probes are possible as well such as neutrinoless double beta decay provided that neutrinos have Majorana mass terms and the light neutrino exchange model is valid. It may also be possible to get directly at the three separate neutrino masses in other ways such as via a precise measurement of the cosmic neutrino background Long et al. (2014) via an experiment such as PTOLEMY Baracchini et al. (2018), although this faces significant experimental challenges for even a detection, let along a separate mass determination Cheipesh et al. (2021); Nussinov and Nussinov (2022); Apponi et al. (2022).
Combining this, our benchmark points [^6] It is important to note that, while our benchmarks assume normal and inverted mass orderings for simplicity, we compute spectra and event rates for all possible orderings and perform a statistical comparison among them. Our results will be similar for different true mass orderings., shown in table 1, are either the minimum masses allowed by oscillations, or the maximum masses allowed by other data sets. We also include one high mass scenario that does not take into account the measured $\Delta m^{2}_{ij}$’s, that could be due to new physics. We have also investigated other scenarios such as the upper limit on the masses from various cosmological data sets including CMB, BAO, and the latest BAO updates from DES and DESI. We found that the constraints in those cases were quite similar to the lowest possible masses.
@@
Figure 6: The expected preferred regions given a detection of a SN at 10 kpc by JUNO under four different hypotheses about the true neutrino masses (see table 1). Note that in this case, only information from the QCD phase transition is used to determine the neutrino masses. The contours are drawn at 1, 2, and 3 $\sigma$ assuming 2 dofs and Wilks’ theorem, and the shaded regions correspond to the 1 $\sigma$ region for 1 dof. The pluses and vertical dashed lines represent the true values.
@@ VII Discussion
The numerical studies have yielded many interesting and nontrivial results. Perhaps the most compelling result found is the fact that in general there is information that can distinguish among the three different mass states in a galactic SN. That is, depending on the true masses of neutrinos, each individual mass state can be uniquely measured. The uniqueness statement is based on the fact that we assume that the PMNS matrix is determined. Thus it is not surprising that the constraints on $m_{3}$ tend to be poorer because $|U_{e3}|$ is much smaller than $|U_{e1}|$ and $|U_{e2}|$. In fact, we believe that any estimate of neutrino masses from a galactic SN should use the full description including the explicit contribution from each separate mass term as described in eqs. 11-14.